Minden Company introduced a new product last year for which it is trying to find an optimal selling price. Marketing studies suggest that the company can increase sales by 5,000 units for each $2 reduction in the selling price. The companys present selling price is $70 per unit, and variable expenses are $40 per unit. Fixed expenses are $540,000 per year. The present annual sales volume (at the $70 selling price) is 15,000 units. Required: 1 What is the present yearly net operating income or loss? Answer Net Operation Loss $90,000 2. What is the present break-even point in unit sales and in dollar sales? Answer: Break- even point units 18,000 Break -even point sales $1260000 NEED Assistance on maximum profit 3. Assuming that the marketing studies are correct, what is the maximum annual profit that the company can earn? At how many units and at what selling price per unit would the company generate this profit?3) Calculate the sales units and selling price at maximum annual profit that the company can earn. Break-even point Break-even point in unit sales x Selling price per unit in dollar sales Therefore, the maximum profit that the company can earn is $270,000, sales unit is 45,000 units and selling price is $58. Working note Calculations are given below Particulars Amount Sales Less: Variable costs Contribution margin Less: Fixed cost Net operating loss S1,050,000 S 600,000 S 450,000 S 540,000 S 90,000 Part 3 Maximum annual profit is $270,000, number of units sold is 45,000 units and selling price is $58. Explanation | Common mistakes | Hint for next step M companys present annual sales are 15,000 units. Present selling price per unit $70, variable cost per unit is $40 and fixed costs are $540,000. Marketing studies suggest that the every $2 reduction in selling price increases the sales units by 5,000 units. Variable cost per unit is remains constant. Contribution margin per unit is calculated by deducting the variable cost per unit from the selling price. Total contribution is calculated by multiplying the sales units with the contribution margin per unit. Net income or loss is calculated by total contribution minus fixed cost. Break-even point in unit sales is calculated by dividing fixed costs by contribution margin per unit. If the selling price is reduced by $2 then the sales units will be increased by 5,000 units. From the above table maximum net income is $270,000 and 45,000 units at $58 selling price per unit would the company to generate this profit.

Can someone please explain where did the 270,000 net income and 45000 units at price 58 came from. I read the explanation but didnt find it helpful at all.

Thank you.
Answer #1

1.Sales = 15,000*70 = $1,050,000

Less: Variable Cost = $600,000

Contribution Margin = $450,000

Fixed Expenses = $540,000

Net Operating Loss = $(90,000)

2.Break even point in units = total fixed costs/contribution margin per unit

= 540,000/30

= 18,000 units

In Dollar Sales = 18,000*70 = $1,260,000

3.Studies say that for every $2 reduction in price, quantity increases by 5,000 units

Selling price per Unit

Units Sold

Variable Cost per Unit

Contribution Margin per Unit

Total Contribution Margin

70

15,000

40

30

450,000

68

20,000

40

28

560,000

66

25,000

40

26

650,000

64

30,000

40

24

720,000

62

35,000

40

22

770,000

60

40,000

40

20

800,000

58

45,000

40

18

810,000

56

50,000

40

16

800,000

Contribution Margin is highest at Selling Price = $58, it starts falling after that.

Hence, best strategy is to sell 45,000 units at $58

Net Income = Contribution Margin – fixed expenses

= $810,000-$540,000

= $270,000